] Archenland has the following tax schedule: the first $20,000 of income is taxed at a rate of 10%; the next $30,000 is taxed at a rate of 25%; beyond this, any additional income is taxed at a rate of 50%. Archenland allows a $4000 exemption per family member.
a] Corin’s family has three members and earns $50,000 per year. Calculate the family’s marginal and average tax rates.
b] Peter’s family has five members and earns $85,000 per year. Calculate the family’s marginal and average tax rates.
Answer #1

Marginal Tax rate is the tax that you would pay on the next dollar of your income. In short it is the Income tax bracket you are in at the moment after deduction of all your exempt earnings.

Average Tax rates is the total taxes you have paid divided by your total taxable income. Average tax rate is generally lesser than the marginal tax rate. The detailed calculation for the above question is worked out as below

Slab ($) Tax Rate
0 - 20000 10%
20001 - 50000 25%
>50000 50%
a) Corin Family b) Peter Family
No of members 3 No of members 5
Exemption per member $4,000 Exemption per member $4,000
Total Exemption 12000 Total Exemption 20000
Income $50,000 Income $85,000
Net taxable income $38,000 Net taxable income $65,000
Tax in Slab 1 =20000*10% = 2000 Tax in Slab 1 =20000*10% = 2000
Tax in Slab 1 $2,000 Tax in Slab 1 $2,000
Tax in Slab 2 =(38000-20000)*25% = 4500 Tax in Slab 2 =(50000-20000)*25% = 7500
Tax in Slab 2 $4,500 Tax in Slab 2 $7,500
Total Income Tax paid $6,500 Tax in Slab 3 =(65000-50000)*50% = 7500
Average Tax =$6500/$38000= 17.10% Tax in Slab 3 $7,500
Marginal Tax rate 25% Total Income Tax paid $17,000
Average Tax =$17000/$65000 =26.15%
Marginal Tax rate 50%